A military plane wants to drop supplies to some survivors on a rocky ndge. plane is flying horizontally at an altitude of 235 m with a speed of 72.5 m/s . How far in advance must the relief goods be released so that they arrive precisely at the survivors' location?
1
Expert's answer
2020-11-11T08:57:10-0500

The initial speed in horizontal direction is "v_{0h} = 72.5m\/s". This speed does not change with time. The initial speed in vertical direction is "v_{0v} = 0m\/s". Thus, the distance travelled in vertial direction is


"h = \dfrac{gt^2}{2}"

where if "h = 235m" then "t" is the time of falling. Thus:


"t = \sqrt{\dfrac{2h}{g}}"

Then the distance travelled in horizontal direction is:


"d = v_{0h}t = v_{0h}\sqrt{\dfrac{2h}{g}}\\
d = 72.5\sqrt{\dfrac{2\cdot 235}{9.81}} \approx 501.8 m"

Answer. The relief goods must be released 501.8 m in advance.